1. Let T: Pn(R) + Pn+1(R) be defined: T(P(x)) = (x + 1)p(x + 2) (a) (2 marks) Show that T is a linear transformation. (b) (3Can someone help in part D AND E PLEASE? solve it in general do not use numbers please
Answer #1

by d) soln Let T Pn (R) → Pnti (IR) be defined by TCP®)) = (x+1) ÞÓx+2) & the derivative operator :D :Pn+1 (R) → Pn (IR) be dcontinuing 1 0 38 2 10 0 3, 0 - =1.2.3..ni m! +0.] p (2) EPnt they are lan find the Corresponding matrix (Dord as an upper tr2007 +6X+4 TOD (x²) (+1) 3(x+2) = 364+1)(x+4x+4) 3<3<3+4x74x + x +4474) - 3(x3 +5x478x+4) ie TOD (x²) 3343 +15x4+24X+]2. So,